Focus (M2) Motion Failure

Friday, December 28, 2012 2:09 PM

Problem(s) Encountered:

Focus was supposed to be at 815 (or perhaps 850) after restart of TCS and

homing of everything the previous day.  However, image of bright star showed telescope very far from focus (very large image of primary instead)

Tried homing focus (home button, select focus drive, select apply). Instead of moving the focus, the focus setting reset from 815 to 0. There was no motion of the focus drive. Tried setting focus to 815.  This moved the focus a long ways, but image was still far out of focus.  Tried moving back to 700.

This command was ignored. Discovered that (1) all commands to set focus to a smaller setting are ignored and (2) the focus home command does not actually home the focus, but resets the focus display to 0 mils.  The command it executes to do this is of the form  home sec_hit=-XXX focus  where XXX is some number like 460.  This was repeatable.  Could not find documentation on focus beyond instructions on commanding it to home using buttons, so out of concern, stopped experimenting and shutdown at 3:30 AM.


We found a broken coupling  from the focus motor shaft to the focus drive mechanism.  After talking to Richard Treffers and found that on a temporary basis that we could make and install a coupling we did. After the focus motor coupler was installed we tested the focus and it work as it should. The current observer was shown what the problem was and the fix so he should be able to get data. There have been couplers ordered which will replace the temp. unit and we will have backups encase this happens again.